# Mitigating Semantic Collapse in Generative Personalization with a Surprisingly Simple Test-Time Embedding Adjustment

### v1

@@ Abstract
In this paper, we investigate the semantic collapsing problem in generative personalization, an under-explored topic where the learned visual concept ($V^{*}$) gradually shifts from its original textual meaning and comes to dominate other concepts in multi-concept input prompts. This issue not only reduces the semantic richness of complex input prompts like "a photo of $V^{*}$ wearing glasses and playing guitar" into simpler, less contextually rich forms such as "a photo of $V^{*}$" but also leads to simplified output images that fail to capture the intended concept. We identify the root cause as unconstrained optimisation, which allows the learned embedding $V^{*}$ to drift arbitrarily in the embedding space, both in direction and magnitude. To address this, we propose a simple yet effective training-free method that adjusts the magnitude and direction of pre-trained embedding at inference time, effectively mitigating the semantic collapsing problem. Our method is broadly applicable across different personalization methods and demonstrates significant improvements in text-image alignment in diverse use cases. Our code is anonymously published at https://anonymous.4open.science/r/Embedding-Adjustment.
@@ 1 Introduction
Text-to-image (T2I) diffusion models have achieved unprecedented fidelity and flexibility in image generation and sparked growing interest in generative personalization [9, 25]. This emerging problem aims to generate images of a specific user-defined visual concept (e.g. a particular person, pet, or object) in different contexts (e.g. on the beach) using a small set of user-provided reference images paired with text prompts describing the desired context. The core objective is to generate visually compelling images that faithfully preserve the unique characteristics of the personal concept while remaining semantically aligned with the textual prompt. Despite recent progress, misalignment between the generated image and the textual prompt is still a major concern.
@@
A robust generative personalization method should allow the user-defined visual concept to be composed with arbitrary contexts in the text prompt without losing fidelity or expressiveness. However, existing approaches often struggle to maintain prompt and generated image alignment, particularly with complex or multi-concept prompts [15, 37]. In these scenarios, the introduced concept can overpower or distort other elements in the prompt, leading to unsatisfactory generations. This issue has commonly been attributed to language drift or overfitting [25], arising from optimising on a limited number of reference images. Beyond overfitting, other factors have also been attributed, such as the limited expressiveness of textual embeddings which compress complex visual concepts into single tokens [36, 20] and the entangled nature of reference sets, where samples may contain co-occurring objects or irrelevant contextual features [1, 14].
@@
In this work, we investigate an under-explored failure mode of generative personalization, which we term the semantic collapsing problem (SCP). SCP refers to the phenomenon where the learned personalized token for a new concept loses its original textual semantic meaning while acquiring increased visual information from the reference concept during the personalization process. As a result, when a user combines the personalized token with other descriptive context in a prompt, the model’s output is overwhelmingly biased towards the personalized concept, often neglecting the other intended elements in the prompt. For example, if one learns a token $V^{*}$ to represent a particular dog, a prompt like “$V^{*}$ playing with a ball in a park” may yield an image that vividly depicts the dog but fails to properly render the ball or the park background. In essence, the prompt’s semantic richness collapses to a simpler form centred on $V^{*}$. While SCP does not severely affect trivial prompts (e.g., “a photo of $V^{*}$”, where the entire prompt is just the concept itself), it undermines the compositionality of the T2I diffusion model on complex prompts. Notably, SCP is distinct from the language drift problem, which describes a different failure where the personalized model overfits to a learned concept and loses generalisation. In SCP, the issue arises when the personalized embedding no longer retains any meaningful textual semantics.
@@
We identify the root cause as unconstrained optimisation, which allows the learned embedding to drift arbitrarily in the embedding space, both in direction and magnitude. We propose a simple yet effective remedy: a training-free, test-time embedding adjustment strategy that realigns the learned concept embedding with its original semantic meaning at inference time. The key idea is to calibrate the embedding’s magnitude and direction to be closer to that of its reference concept. This adjustment is done without altering the model weights or requiring any additional training. The embedding is modified on-the-fly before image generation. By enforcing a small rotation and rescaling in the text encoder’s latent space, we constrain the personalized token to behave more like a regular word, ensuring that it contributes to the image generation in balance with other tokens in the prompt. Our approach is lightweight and broadly compatible with different personalization methods and demonstrates significant improvements in text-image alignment in diverse use cases.
@@
In summary, our contributions are as follows: We define the semantic collapsing problem (SCP) in generative personalization problem, and provide an empirical analysis of its existence in both textual and image spaces. Our analysis reveals its root cause, which is the unconstrained optimisation. We propose the test-time embedding adjustment, a novel solution that requires no additional training, to mitigate SCP by aligning the learned embedding’s direction and norm with its original semantic concept. We demonstrate that our proposed approach can be applied into different personalization methods such as Textual Inversion [9], DreamBooth [25], and Custom Diffusion [16] to significantly improve image generation in complex prompts across a wide range of scenarios.
(a) Intermediate generated images
(b) Alignment scores
Figure 1: Illustration of the SCP on Textual Inversion. Left: The intermediate generated images $\hat{x}$ of a prompt ‘a photo of a $V^{*}$ wearing glasses and writing in a red notebook’. The generated image is gradually biased towards the personalized concept $V^{*}$ (i.e., easier to recognize as ‘Henry Cavill’) and loses the context $p$ (i.e., harder to recognize as ‘writing in a red notebook’) through out the personalization process. Right: The alignment scores (average over 100 random seeds) which empirically validate the SCP. The alignment $S(\hat{x},p)$ ($\square$) with the context $p$ drops over time, while the alignment $S(\hat{x},x_{gt})$ ($+$) with the ground truth $x_{gt}$ increases.
2 Personalizing T2I Generative Models
Given a text-to-image diffusion model $\epsilon_{\theta}$, where $\epsilon_{\theta}(x_{t},t,p)$ represents the predicted noise at time step $t$ given the textual embedding $\tau(p)$ of a prompt $p$ and the noisy intermediate vector $x_{t}$ [11, 29, 24]. And given a set of personal images $\mathcal{X}=\{x_{1},x_{2},\cdots,x_{n}\}$ and a pre-trained T2I model $\epsilon_{\theta}$, the goal of generative personalization is to identify a textual embedding $v^{*}$ associated with a specific verbalizable keyword $V^{*}$ (e.g., ‘sks’, ‘<new>’, etc.). This keyword represents the implicit visual concept shared in the reference set $\mathcal{X}$, enabling the model to generate images with the personal concept using any textual prompt $p$ containing the keyword $V^{*}$, e.g., $\lfloor p,V^{*}\rfloor=$ ’A photo of $V^{*}$ playing on a beach’, where .,$\lfloor$.,$\rfloor$ . , . ⌋ is the sentence construction operator.
This keyword represents the implicit visual concept shared in the reference set $\mathcal{X}$, enabling the model to generate images with the personal concept using any textual prompt $p$ containing the keyword $V^{*}$, e.g., $\lfloor p,V^{*}\rfloor=$ ’A photo of $V^{*}$ playing on a beach’, where .,$\lfloor$.,$\rfloor$ . , . ⌋ is the sentence construction operator. We denote $M$ is the embedding matrix of the entire vocabulary of the text encoder $\tau$, and $M_{k}$ is a specific row of the matrix corresponding to the token $k\in\text{vocab}_{\tau}$, where $k$ can be $V^{*}$ (i.e., $v^{*}=M_{V^{*}}$) or any arbitrary token like ‘dog’, ‘cat’, etc. We denote $\hat{x}=G(p)$ as the image generation model that takes a prompt $p$ as input and outputs an image $\hat{x}$. We discuss two representative methods for personalizing T2I generative models: Textual Inversion (TI) [9] and DreamBooth [25] and other related works in Appendix A.
@@ 3.1 Semantic Collapsing Problem
Comparison with Other Challenges in Generative personalization. First, we emphasise that SCP is not specific to the two representative methods we study (TI and DB), but is a general issue in personalization. Second, SCP is distinct from other recognised challenges (ref. Section A) such as the language drift problem [25], which describes how the personalized model $\epsilon_{\theta}$ overfits to a learned concept and loses generalisation, for instance, generating a personalized dog image from a prompt like ‘a photo of a dog’, even when no personalized keyword is present. In contrast, SCP refers to the case where the keyword $V^{*}$ no longer retains any meaningful textual semantics, instead encoding only visual information. Third, while SCP contributes to the broader challenge of misalignment between generated images and prompts, a major concern in generative personalization, it stems from a specific cause: the unconstrained optimisation of the embedding during personalization, which has not been thoroughly studied in prior work.
@@ 3.2 Empirical Hunting for SCP
In this section, we present empirical evidence supporting the existence of the semantic collapsing problem and its impact on generation quality. Our key findings are as follows: Existence of SCP. SCP exists in the textual domain, where the prompt $\lfloor p,V^{*}\rfloor$ is dominated by the learned embedding $V^{*}$ and the semantic meaning of the entire prompt gradually collapses to the learned embedding $V^{*}$, i.e., $\tau(\lfloor p,V^{*}\rfloor)\rightarrow\tau(V^{*})$. Negative Impact on Generation Quality. SCP leads to the degradation/misalignment in generation quality in the image space, i.e., $G(\lfloor p,V^{*}\rfloor)\rightarrow G(V^{*})$, particularly for prompts with complex semantic structures. Surprisingly Positive Impact. SCP can also lead to the positive impact on generation quality, particularly for prompts where the concept $c$ requires a strong visual presence to be recognisable. Root Cause of SCP. SCP arises from unconstrained optimisation during personalization, which leads to arbitrary shifts (both in magnitude and direction) in the embedding of $V^{*}$ away from its original semantic concept $c$.
Figure 2: (a) The inter-set distance $d(P_{V^{*}},P_{c})$ and intra-set distance $d(P_{V^{*}},P_{V^{*}})$ over the personalization process, and (b) The distance between all possible pairs of sets, notably $d(P_{V^{*}},P_{V^{*}}^{\text{simple}})$.
@@ 3.2.1 Empirical Evidence for SCP in Textual Space
Recall our hypothesis: a keyword $V^{*}$ initialised from a concept $c$ to capture a visual target $v_{gt}$ will lose its semantic meaning and dominate any arbitrary context $p$ when combined into a prompt. A straightforward way to evaluate this semantic shift is to compare the embedding vectors $M_{V^{*}}$ and $M_{c}$. However, modern LLMs and Diffusion Models use contextualised text embeddings, in which the surrounding context significantly influences the final representation of a prompt.
To address this, we propose measuring the difference between $V^{*}$ and $c$ in the presence of a diverse set of contextual prompts $A=\{a_{1},a_{2},\cdots,a_{n}\}$. These are generated by querying an LLM with the instruction: "Write 200 sentences with diverse topics and contents. Each sentence should be 10–30 words long and must include the keyword $c$." Sample sentences are provided in Table 3, and the full dataset and prompts can be found in the anonymous repository.
We then construct two sets of prompts: $P_{V^{*}}=\{\lfloor a_{i},V^{*}\rfloor\}_{a_{i}\in A}$ and $P_{c}=\{\lfloor a_{i},c\rfloor\}_{a_{i}\in A}$, which are used to assess the contextualised difference between $V^{*}$ and $c$.
[…]
@@ Results.
Figure 2(a) shows the average inter-set and intra-set distances, $d(P_{V^{*}},P_{c})$ and $d(P_{V^{*}},P_{V^{*}})$, measured over training iterations. It can be seen that the inter-set distance $d(P_{V^{*}},P_{c})$ increases steadily over time, indicating that the learned embedding $v^{*}$ progressively diverges from its initial textual semantic meaning $c$.
Interestingly, the intra-set distance $d(P_{V^{*}},P_{V^{*}})$ decreases over time, suggesting that the embeddings of prompts $\lfloor a_{i},V^{*}\rfloor$ within $P_{V^{*}}$ become less diverse and more similar to one another. This reflects a growing dominance of the learned embedding $v^{*}$ across prompts, effectively overriding the contextual variations in $a_{i}$ and becoming the principal semantic component of each prompt.
@@ 3.2.2 The Two-Way Impacts on Personalization
In the previous section, we demonstrated the semantic collapsing problem in the textual space, where the prompt $\lfloor p,V^{*}\rfloor$ becomes increasingly dominated by the learned embedding $V^{*}$, causing the semantic meaning of the entire prompt to collapse toward $V^{*}$. In this section, we extend our analysis to investigate how this phenomenon impacts image generation quality.
Specifically, we generate personalized images $\hat{x}=G(\lfloor p,V^{*}\rfloor)$ using a list of prompts $P$ (e.g., ‘a photo of a $V^{*}$ man holding a cat’), with 100 images generated per prompt. We evaluate the generation quality using the CLIP-Image-Image alignment score $S_{I}=S(\hat{x},x_{gt})$, which measures the similarity between the generated image $\hat{x}$ and the ground-truth image $x_{gt}$ of the reference concept. In addition, we compute three CLIP-Text-Image alignment scores: (i) $\text{CLIP}_{T}^{p}$ or $S_{T}^{p}=S(\hat{x},p)$ alignment with the contextual part $p$ (e.g., ‘holding a cat’), and (ii) $\text{CLIP}_{T}^{c}$ or $S_{T}^{c}=S(\hat{x},c)$ alignment with the original concept $c$ (e.g., ‘a man’). (iii) $\text{CLIP}_{T}^{f}$ or $S_{T}^{f}=S(\hat{x},\lfloor p,c\rfloor)$ alignment with the full prompt.
@@
(Unsurprising) The image-to-image alignment score $S_{I}$ increases over time ($+$ line), indicating that $V^{*}$ effectively captures the visual appearance of the target concept. This confirms that the learned embedding $V^{*}$ successfully personalizes the visual identity from the reference set.
@@
[…]

### v2

# Mitigating Semantic Collapsing Problem in Generative Personalization with Test-time Embedding Adjustment

@@ Abstract
In this paper, we investigate the semantic collapsing problem in generative personalization, an under-explored topic where the learned visual concept ($V^{*}$) gradually shifts from its original textual meaning and comes to dominate other concepts in multi-concept input prompts. This issue not only reduces the semantic richness of complex input prompts like "a photo of $V^{*}$ wearing glasses and playing guitar" into simpler, less contextually rich forms such as "a photo of $V^{*}$" but also leads to simplified output images that fail to capture the intended concept. We identify the root cause as unconstrained optimisation, which allows the learned embedding $V^{*}$ to drift arbitrarily in the embedding space, both in direction and magnitude. To address this, we propose a simple yet effective training-free method that adjusts the magnitude and direction of pre-trained embedding at inference time, effectively mitigating the semantic collapsing problem. Our method is broadly applicable across different personalization methods and demonstrates significant improvements in text-image alignment in diverse use cases. Our code is published at https://github.com/tuananhbui89/Embedding-Adjustment.
@@ 1 Introduction
Text-to-image (T2I) diffusion models have achieved unprecedented fidelity and flexibility in image generation and sparked growing interest in generative personalization (Gal et al., 2022; Ruiz et al., 2023). This emerging problem aims to generate images of a specific user-defined visual concept (e.g. a particular person, pet, or object) in different contexts (e.g. on the beach) using a small set of user-provided reference images paired with text prompts describing the desired context. The core objective is to generate visually compelling images that faithfully preserve the unique characteristics of the personal concept while remaining semantically aligned with the textual prompt. Despite recent progress, misalignment between the generated image and the textual prompt is still a major concern.
@@
A robust generative personalization method should allow the user-defined visual concept to be composed with arbitrary contexts in the text prompt without losing fidelity or expressiveness. However, existing approaches often struggle to maintain prompt and generated image alignment, particularly with complex or multi-concept prompts (Kong et al., 2024; Zhu et al., 2025). As illustrated in Figure˜1, the user-defined concept can overpower or distort other elements in the prompt, leading to unsatisfactory generations. This issue has commonly been attributed to language drift - a phenomenon where the model gradually forgets how to generate its pretrained concepts and instead becomes overly focused on the user-defined ones (Lee et al., 2019). This drift typically stems from overfitting on a limited number of reference images (Ruiz et al., 2023). Beyond overfitting, other factors have also been attributed, such as the limited expressiveness of textual embeddings which compress complex visual concepts into single tokens (Zhang et al., 2023; Mou et al., 2024) and the entangled nature of reference sets, where samples may contain co-occurring objects or irrelevant contextual features (Avrahami et al., 2023; Jin et al., 2024). While misalignment between textual prompt and generated images is a well-recognized challenge in generative personalization, with numerous mitigation strategies proposed, including latent optimization (Rassin et al., 2023), regularization (Han et al., 2023; Qiu et al., 2023; Arar et al., 2024) and concept disentanglement (Motamed et al., 2024; Huang et al., 2024a), the underlying causes and mechanisms remain relatively underexplored.
Figure 1: Our Test-time Embedding Adjustment (TEA) method consistently enhances text-image alignment across diverse personalization approaches (Textual Inversion, DreamBooth, and their variants) and architectures (Stable Diffusion, Flux). Notably, TEA also counteracts the anti-personalization effect of Anti-DreamBooth and restores the protected concept.
@@
To gain deeper insight into the misalignment problem, this paper presents an empirical investigation into the dynamics of learned personalised tokens throughout the personalisation process. Our analysis reveals an intriguing phenomenon: the personalised token gradually loses its original textual semantic meaning while acquiring increased visual information from the reference images. When such a token is used in a prompt with rich descriptive text, the generated image becomes disproportionately semantically dominated by the personalized concept, often neglecting the other intended elements in the prompt. For example, in Figure˜1, if one learns a token $V^{*}$ to represent a particular cat, a prompt like $V^{*}$ in grand canyon” may yield an image that vividly depicts the cat but fails to properly render the grand canyon background. Essentially, the prompt’s semantic complexity collapses to a simplified form centred on $V^{*}$. We refer to this phenomenon as the semantic collapsing problem (SCP). Unlike language drift where the personalized model overfits to a learned concept and loses its ability to generate other pretrained concepts, SCP arises when the personalized embedding collapses, no longer retaining meaningful textual semantics and instead encoding primarily the visual information of the reference concept. While SCP might not severely affect trivial prompts (e.g., “a photo of $V^{*}$”), it undermines the compositionality of the T2I diffusion model on complex prompts.
@@
We identify the root cause of SCP as unconstrained optimisation, which allows the learned embedding to drift arbitrarily in the embedding space, both in direction and magnitude. We propose a simple yet effective remedy: a training-free, test-time embedding adjustment (TEA) strategy that realigns the learned concept embedding with its original semantic meaning at inference time. The key idea is to calibrate the embedding’s magnitude and direction to be closer to that of its reference concept. This adjustment is done without altering the model weights or requiring any additional training. The embedding is modified on-the-fly before image generation. By enforcing a small rotation and rescaling in the text encoder’s latent space, we constrain the personalized token to behave more like a regular word, ensuring that it contributes to the image generation in balance with other tokens in the prompt. Our approach is lightweight and broadly compatible with almost all existing personalization methods and demonstrates significant improvements in text-image alignment in diverse use cases. Surprisingly, beyond improving personalization quality, we also uncover a surprising vulnerability in anti-personalization frameworks like Anti-DreamBooth (Van Le et al., 2023) under the lens of SCP. In particular, when applying TEA to models poisoned by Anti-DreamBooth, we find that TEA can partially reverse adversarial corruption and restore more faithful generations of the protected concept (see Figure˜1). This result highlights a false sense of security in current anti-personalization defenses and provides new insights into their limitations.
@@
In summary, our contributions are as follows: We define the semantic collapsing problem (SCP) in generative personalization problem, and provide an empirical analysis of its existence in both textual and image spaces. Our analysis reveals its root cause, which is the unconstrained optimisation. We propose the test-time embedding adjustment, a novel solution that requires no additional training, to mitigate SCP by aligning the learned embedding’s direction and norm with its original semantic concept. We demonstrate that our proposed approach can be applied into almost all existing personalization frameworks such as Textual Inversion (Gal et al., 2022), DreamBooth (Ruiz et al., 2023), Custom Diffusion (Kumari et al., 2023), EasyControl (Zhang et al., 2025), ReVersion (Huang et al., 2024b), and ClassDiffusion (Huang et al., 2024a) to significantly improve image generation in complex prompts across a wide range of scenarios. We show that TEA unexpectedly mitigates adversarial corruption introduced by Anti-DreamBooth, revealing an overlooked weakness in anti-personalization defences.
@@ 2 Semantic Collapsing Problem in Generative Personalization
2.1 Terminologies
Given a set of personal images $\mathcal{X}$ and a pre-trained T2I model $\epsilon_{\theta}$, the goal of generative personalization is to identify a textual embedding $v^{*}$ associated with a specific verbalizable keyword $V^{*}$ (e.g., ‘sks’, ‘<new>’, etc.). This keyword represents the implicit visual concept shared in the reference set $\mathcal{X}$, enabling the model to generate images with the personal concept using any textual prompt $p$ containing the keyword $V^{*}$, e.g., $\lfloor p,V^{*}\rfloor=$ ’A photo of $V^{*}$ playing on a beach’, where $\lfloor.,.\rfloor$ is the sentence construction operator. We denote $c$ as the reference semantic concept of the keyword $V^{*}$.
We denote $M$ is the embedding matrix of the entire vocabulary of the text encoder $\tau$, and $M_{k}$ is a specific row of the matrix corresponding to the token $k\in\text{vocab}_{\tau}$, where $k$ can be $V^{*}$ (i.e., $v^{*}=M_{V^{*}}$) or any arbitrary token like ‘dog’, ‘cat’, etc. We denote $\hat{x}=G(p)$ as the image generation model that takes a prompt $p$ as input and outputs an image $\hat{x}$.
@@ 2.2 Semantic Collapsing Problem
Comparison with Other Challenges in Generative personalization. First, we emphasise that SCP is not specific to the two representative methods we study (TI and DB), but is a general issue in personalization. Second, SCP is distinct from other recognised challenges (ref. Section A) such as the language drift problem (Ruiz et al., 2023), which describes how the personalized model $\epsilon_{\theta}$ overfits to a learned concept and loses generalisation. Third, while SCP contributes to the broader challenge of misalignment between generated images and prompts, a major concern in generative personalization, it stems from a specific cause: the unconstrained optimisation of the embedding during personalization, which has not been thoroughly studied in prior work.
@@ 2.3 Empirical Hunting for SCP
In this section, we present empirical evidence supporting the existence of the semantic collapsing problem and its impact on generation quality. Our key findings are as follows:
❶Existence of SCP. SCP exists in the textual domain, where the prompt $\lfloor p,V^{*}\rfloor$ is dominated by the learned embedding $V^{*}$ and the semantic meaning of the entire prompt gradually collapses to the learned embedding $V^{*}$, i.e., $\tau(\lfloor p,V^{*}\rfloor)\rightarrow\tau(V^{*})$.
❷Negative Impact on Generation Quality. SCP leads to the degradation/misalignment in generation quality in the image space, i.e., $G(\lfloor p,V^{*}\rfloor)\rightarrow G(V^{*})$, particularly for prompts with complex semantic structures.
❸Surprisingly Positive Impact. SCP can also lead to the positive impact on generation quality, particularly for prompts where the concept $c$ requires a strong visual presence to be recognisable.
❹Root Cause of SCP. SCP arises from unconstrained optimisation during personalization, which leads to arbitrary shifts (both in magnitude and direction) in the embedding of $V^{*}$ away from its original semantic concept $c$.
@@ 2.3.1 Empirical Evidence for SCP in Textual Space
Recall our hypothesis: a keyword $V^{*}$ initialised from a concept $c$ to capture a visual target $v_{gt}$ will lose its semantic meaning and dominate any arbitrary context $p$ when combined into a prompt.
To verify this, we propose measuring the difference between $V^{*}$ and $c$ in the presence of a diverse set of contextual prompts $A=\{a_{1},a_{2},\cdots,a_{n}\}$. These are generated by querying an LLM with the instruction: "Write 200 sentences with diverse topics and contents. Each sentence should be 10–30 words long and must include the keyword cc." Sample sentences are provided in Table 3. We then construct two sets of prompts: $P_{V^{*}}=\{\lfloor a_{i},V^{*}\rfloor\}_{a_{i}\in A}$ and $P_{c}=\{\lfloor a_{i},c\rfloor\}_{a_{i}\in A}$, which are used to assess the contextualised difference between $V^{*}$ and $c$.
To quantify this difference, we compute distances between the sets $P_{V^{*}}$ and $P_{c}$ using four metrics: Euclidean, Hausdorff, Mahalanobis, and KL divergence. We also evaluate intra-set variability $d(P_{V^{*}},P_{V^{*}})$ measuring the separation among items within $P_{V^{*}}$. The distance metrics are summarised in Table 5. We use Textual Inversion (TI) and DreamBooth (DB) to learn a personalized human face concept. The training data comprises 16 images from the CelebA dataset (Liu et al., 2015) (subject ID: 342, ‘Henry Cavill’). The embedding $v^{*}$ is initialised using the concept $c=\text{`man'}$ for TI, and $c=\text{`a photo of a sks man'}$ for DB.
Figure 2: (a/left) The inter-set distance $d(P_{V^{*}},P_{c})$ and intra-set distance $d(P_{V^{*}},P_{V^{*}})$ over the personalization process, and (b) The distance between all possible pairs of sets, notably $d(P_{V^{*}},P_{V^{*}}^{\text{simple}})$.
@@ Results.
Figure 2(a) shows the average inter-set and intra-set distances, $d(P_{V^{*}},P_{c})$ and $d(P_{V^{*}},P_{V^{*}})$, measured over training iterations. It can be seen that the inter-set distance $d(P_{V^{*}},P_{c})$ increases steadily over time, indicating that the learned embedding $v^{*}$ progressively diverges from its initial textual semantic meaning $c$. Interestingly, the intra-set distance $d(P_{V^{*}},P_{V^{*}})$ decreases over time, suggesting that the embeddings of prompts $\lfloor a_{i},V^{*}\rfloor$ within $P_{V^{*}}$ become less diverse and more similar to one another. This reflects a growing dominance of the learned embedding $v^{*}$ across prompts, effectively overriding the contextual variations in $a_{i}$ and becoming the principal semantic component of each prompt.
@@ 2.3.2 The Two-Way Impacts on Personalization
In this section, we extend our analysis to investigate how SCP impacts image generation quality. Specifically, we generate personalized images $\hat{x}=G(\lfloor p,V^{*}\rfloor)$ using a list of prompts $P$ (e.g., ‘a photo of a $V^{*}$ man holding a cat’), with 100 images generated per prompt. We evaluate the generation quality using the CLIP-Image-Image alignment score $S_{I}=S(\hat{x},x_{gt})$, which measures the similarity between the generated image $\hat{x}$ and the ground-truth image $x_{gt}$ of the reference concept. In addition, we compute three CLIP-Text-Image alignment scores: (i) $\text{CLIP}_{T}^{p}$ or $S_{T}^{p}=S(\hat{x},p)$ alignment with the contextual part $p$ (e.g., ‘holding a cat’), and (ii) $\text{CLIP}_{T}^{c}$ or $S_{T}^{c}=S(\hat{x},c)$ alignment with the original concept $c$ (e.g., ‘a man’). (iii) $\text{CLIP}_{T}^{f}$ or $S_{T}^{f}=S(\hat{x},\lfloor p,c\rfloor)$ alignment with the full prompt.
@@
(Unsurprising) The image-to-image alignment score $S_{I}$ increases over time (++ line), indicating that $V^{*}$ effectively captures the visual appearance of the target concept. This confirms that the learned embedding $V^{*}$ successfully personalizes the visual identity from the reference set.
@@
(Unexpected Positive Effect) For some prompts, the concept-to-image alignment score $S_{T}^{c}$ actually increases over time ($\lozenge$ line). Early in training, prompts with $c$ (e.g., man) often generate images dominated by the context $p$. As training progresses, the learned embedding $V^{*}$ reduces this context dominance, strengthening the representation of $c$. This effect is most evident when $c$ demands a strong visual presence (e.g., man in “man writing in a red notebook”), where close-up subject renderings naturally downplay the surrounding context.